Dear sir/madam, please help me to answer this...
Trial1 Trial2 Trial3
Mass of bottle+KhP 98.108g 109.097g 112.442g
Mass of bottle 97.526g 108.51g 111.89g
Mass of KHP used 0.5819g 0.5868g 0.5523g
Final buret reading 32.4ml 29.4ml 24.3ml
Initial buret reading 0 0 0
mL of NaOH used 32.4ml 29.4ml 24.3ml
Molarity of NaOH ________ ________ ________

Average molarity_____________
Standard deviation ____________
1
Expert's answer
2019-12-16T04:53:55-0500

Molar mass of KHP=204.22mol/L

number of moles(KHP)=m/M

trial1) 0.5819/204.22=0.00285

trial2) 0.5868 /204.22=0.00287

trial3) 0.5523/204.22=0.00270g

number of moles(KHP)=number of moles(NaOH)

Molarity(NaOH)=n/V

trial1)0.00285mol/0.324L=0.0088mol/L

trial2)0.00287mol/0.294L=0.0098mol/L

trial3)0.0027mol/0.243L=0.0111mol/L

Average molarity=(0.0088+0.0098+0.0111)/3=0.0099mol/L

Standard deviation:

(0.0088-0.0099)2=0.00000121

(0.0098-0.0099)2=0.00000001

(0.0111-0.0099)2=0.00000144

Standard derivation2=(0.00000121+0.00000001+0.00000144)/3=0.00000089

standard derivation= 0.00094